OpenStudy (anonymous):

There are two numbers. When one is added to each of the two numbers, their ratio is 1 : 2. When 5 is subtracted from each of them, their ratio is 5 : 11. Find the numbers.

OpenStudy (anonymous):

x+1/y+1=1/2 so 2x+1=y+1 so y=2x

OpenStudy (anonymous):

x-5/2x-5=5/11, 11x-55=10x-25 so x=30

OpenStudy (anonymous):

so y =2x then y=60?

OpenStudy (dannyrod2000):

i don't know how to explain this to you but good effort on your work :)

OpenStudy (anonymous):

no no...\[\frac{ x+1 }{ y+1 } = \frac{ 1 }{ 2 } \Longrightarrow 2x + 2 = y + 1\]

OpenStudy (anonymous):

oh yes.i made a mistake there .3q @PFEH.1999

OpenStudy (anonymous):

and so try it again...

OpenStudy (anonymous):

so 35 n 71

OpenStudy (anonymous):

i will always make mistake by not multiplying other one.

OpenStudy (anonymous):

oh no problem :) if u practice a lot you won't make mistakes :)

OpenStudy (anonymous):

and so u should notice this: \[\frac{ x-5 }{ y-5 } = \frac{ 5 }{ 11 }\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

35 n 71

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

above i gave the answer

OpenStudy (anonymous):

doesn't seem correct ... let me check it again.

OpenStudy (anonymous):

maybe i made a mistake.

OpenStudy (anonymous):

correct ;) x=35 and y=71 i made a mistake at first ;)

OpenStudy (anonymous):

see=> y=2x+1.so x-5/2x+1-5=5/11

OpenStudy (anonymous):

same as me

OpenStudy (anonymous):

ok then ;)

OpenStudy (anonymous):

do u still have any problems?

OpenStudy (anonymous):

yes .i will post if i have doubt.3q @PFEH.1999 .i need trigonometry notes.u gave formula for tan(a+b)=tana*tanb/1-tana*tan b.thats simple but i dont understand c2,c1,b mentioning in triangle for the sum tan(-15)

OpenStudy (anonymous):